What value of r makes -6r+2(r+2)-5=-20 true?

Answer by tommyt3rd(5050)

-6r+2(r+2)-5=-20
-6r+2r+4-5=-20
-4r-1=-20
-4r=-20+1
-4r=-19

r=19/4



:)